OpenStudy (anonymous):

help can someone help me simplify y^2-25/5-y

OpenStudy (anonymous):

\[y^2 - 5 - y = y^2 - y - 5 \]

OpenStudy (anonymous):

\[{y^2-25 \over 5-y}={(y-5)(y+5) \over 5-y}=-(y+5)\]

OpenStudy (anonymous):

this is by simplifying by removing factpors of 1